# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

### v2

@@ D.4 Impact on Input Prediction
GPT-4 Omni GPT-4o-Mini $57.67$ $69.52$ $-17.21$ $-13.65$ $+0.75$ $+0.73$ $+1.00$ $+0.73$ $-17.58$ $-6.02$
GPT-4o $68.04$ $79.12$ $-9.04$ $-5.88$ $+3.83$ $+1.00$ $+0.96$ $+2.12$ $-5.67$ $-2.25$
Claude Series Claude-3.5-Haiku-20241022 $54.67$ $63.62$ $-9.29$ $-10.62$ $+3.83$ $+2.62$ $+4.25$ $+1.50$ $-24.79$ $-19.00$
Claude-3.5-Sonnet-20241022 $73.33$ $80.12$ $-11.67$ $-7.75$ $+0.08$ $-2.75$ $-1.25$ $-2.12$ $-18.21$ $-13.62$
Gemini Series Gemini-1.5-Flash $57.62$ $73.62$ $-17.38$ $-17.50$ $+5.88$ $-4.00$ $+3.46$ $-3.12$ $-17.83$ $-8.12$
Gemini-2.0-Flash $70.50$ $84.88$ $-29.12$ $-11.38$ $-0.54$ $-2.38$ $-8.08$ $-10.25$ $-15.17$ $-3.62$
Gemini-1.5-Pro-002 $70.87$ $81.88$ $-14.79$ $-12.25$ $-1.75$ $-5.75$ $-2.67$ $-7.88$ $-13.25$ $-6.25$
DeepSeek DeepSeek-V3 $69.08$ $82.88$ $-11.46$ $-6.62$ $+0.77$ $-0.25$ $+1.08$ $+0.75$ $-3.88$ $+2.62$
LLaMA Series LLaMA-3.1-8B-Instruct $37.12$ $41.62$ $-11.33$ $-25.75$ $+1.44$ $+0.62$ $+1.13$ $-1.50$ $-2.67$ $-10.38$
LLaMA-3.1-70B-Instruct $62.00$ $66.38$ $-17.54$ $-11.87$ $+3.04$ $+3.00$ $+1.00$ $+1.12$ $-10.79$ $-6.75$
LLaMA-3.1-405B-Instruct $66.79$ $75.00$ $-10.21$ $-6.75$ $+3.92$ $+2.75$ $+2.00$ $+1.75$ $-5.21$ $-4.25$
LLaMA-3.3-70B-Instruct $63.38$ $76.50$ $-18.62$ $-12.00$ $+0.79$ $-0.25$ $-2.58$ $+0.50$ $-13.83$ $-8.50$
Qwen Series Qwen2.5-7B-Instruct $38.83$ $51.38$ $-17.25$ $-25.88$ $+5.71$ $+1.62$ $+4.75$ $+3.87$ $-3.63$ $-7.38$
Qwen2.5-14B-Instruct $50.54$ $60.75$ $-19.42$ $-9.13$ $+4.38$ $+6.62$ $+2.46$ $+3.12$ $-4.42$ $-0.75$
Qwen2.5-32B-Instruct $63.50$ $74.12$ $-16.54$ $-15.12$ $+4.92$ $+0.38$ $+1.42$ $-3.00$ $-13.08$ $-8.12$
Qwen2.5-72B-Instruct $64.46$ $73.81$ $-17.31$ $-7.29$ $+1.71$ $+1.19$ $-2.38$ $+1.81$ $-16.46$ $-5.06$
Qwen2.5-Coder-32B-Instruct $74.25$ $78.25$ $-21.21$ $-16.12$ $-0.75$ 0.00 $-1.79$ $+0.25$ $-27.08$ $-26.12$
Average $\mathbf{-15.85}$ $\mathbf{-12.68}$ $\mathbf{+2.24}$ $\mathbf{+0.30}$ $\mathbf{+0.28}$ $\mathbf{-0.61}$ $\mathbf{-12.56}$ $\mathbf{-7.86}$
Appendix E Case Studies
E.1 Case Study 1: How does DeepSeek-V3 be misled by ALL
@@
Code 19: ALL-perturbed example from Crux (sample 528) for output prediction.
@@ 18 Var_6 = Var_6 + Var_3
E.2 Case Study 2: How does GPT-4o be misled by MDC
@@
Code 22: MDC-perturbed example from Crux (sample 1) for output prediction.
@@
E.3 Case Study 3: How does Qwen2.5-72B-Ins be misled by MPS
This case (Code 24) highlights how MPS can subtly manipulate model reasoning. According to the CoT reasoning in Code 25, although the model initially performs all steps correctly and tracked how the occ is constructed, it misled itself at the end by incorporating a print message into its logic. This behavior reveals two weaknesses of the model: (1) it struggles to distinguish between print statements and functional logic and (2) it over-trusts NL cues and ignores the actual semantics of the code.
@@ 5. For each character ch in "URW rNB":
E.4 Case Study 4: How does MHC override Claude-3.5-Sonnet’s answer
@@ 5. Final value of ans is 4
E.5 Case Study 5: How does GPT-4o gain from reasoning shortcuts
@@
This function applies a sequence of operations to the input string text: (1) @ reverses the string, (2) ~ converts it to uppercase, and (3) any other character removes itself from the end of the string if it matches the currently last character. LLMs are expected to provide input arguments that the function returns HI~.
In the vanilla thought, as shown in Code 30, GPT-4o mentions, “Text is hi~, ~ makes it HI~, @ reverses it to ~IH ”, it wrongly assumes ~IH is the final output because it applies the reversal after the uppercasing step, resulting in an incorrect answer.
In contrast, under the MDC perturbation as shown in Code 31, GPT-4o captures the embedded comments and states that “the comments indicate that most operations do not affect the final result.” Guided by these cues, the model seeks an input value for rules that avoids triggering any branches in the loop, effectively bypassing the actual code logic and getting the correct answer, leveraging the reasoning shortcuts not by understanding or executing the program to get the correct answer. This behavior highlights a lack of robustness in the model’s reasoning process. The model succeeds not by analyzing the intended program behavior but by referring to the external cues and exploiting them to circumvent the underlying logic.
@@
Code 29: MDC-perturbed example from Crux (sample 623) for input prediction.
@@
E.6 Case Study 6: How does QwQ-32B Collapse
In §5.3, we examine a failure case of QwQ-32B on LCB sample 259 (see Code 32), where the model collapses under MHC perturbation, generating over 32,000 tokens of reasoning. We manually segmented the reasoning process into six stages (see Figure 4):
@@ 3 A = deque()
Figure 4: Breakdown of QwQ-32B’s reasoning trajectory under MHC perturbation on LCB sample 259. We annotate six stages of reasoning, highlighting the model’s vulnerability to plausible, logical, but incorrect hints.
Appendix F Limitation
This study is subject to several limitations. Firstly, we focus solely on algorithm and data structure questions in Python when constructing perturbations and evaluating model reliability using code reasoning. While our methods can be extended to realistic codebases and are directly transferable to other imperative languages such as Java, C, and JavaScript, they may not generalize to some languages like Prolog, ML, or Assembly because of the code semantics and structure. Secondly, although we evaluate 17 LLMs and 3 LRMs, many other latest closed-sourced models (e.g., GPT-4.5, o3, o4-Mini, Gemini-2.5, and Claude-3.7) and proprietary open-sourced models (e.g., Mixtral, Falcon, Vicuna) are not included due to budget constraints. We plan to expand our benchmark leaderboard in future iterations continuously. Thirdly, we adopt pass@1 as the primary evaluation metric. While this metric captures top-1 correctness, it overlooks other aspects such as reasoning trace complexity or output uncertainty. Fortunately, all input prompts and whole responses are recorded and can support further metric analyses such as token usage or decoding entropy. Lastly, our work does not offer definitive mitigation strategies. Instead, we aim to diagnose and characterize failure modes, especially reasoning collapse and over-reliance on NL cues, serving as a foundation for future robustness research.

### v3

@@ E.3 Additional Performance Degradation Tables for Input Prediction
GPT-4 Omni GPT-4o-Mini $57.8$ $69.5$ 17.3-17.3 13.6-13.6 +0.7+0.7 +0.7+0.7 +0.8+0.8 +0.7+0.7 17.7-17.7 6.0-6.0
GPT-4o $68.0$ $79.1$ 9.0-9.0 5.9-5.9 +3.8+3.8 +1.0+1.0 +1.0+1.0 +2.1+2.1 5.7-5.7 2.2-2.2
Claude Series Claude-3.5-Haiku-20241022 $54.7$ $63.6$ 9.3-9.3 10.6-10.6 +3.7+3.7 +2.6+2.6 +4.2+4.2 +1.5+1.5 24.8-24.8 19.0-19.0
Claude-3.5-Sonnet-20241022 $73.3$ $80.1$ 11.7-11.7 7.8-7.8 +0.1+0.1 2.8-2.8 1.3-1.3 2.1-2.1 18.2-18.2 13.6-13.6
Gemini Series Gemini-1.5-Flash $57.6$ $73.6$ 17.2-17.2 17.5-17.5 +5.9+5.9 4.0-4.0 +3.5+3.5 3.1-3.1 17.8-17.8 8.1-8.1
Gemini-2.0-Flash $70.5$ $84.9$ 29.7-29.7 11.4-11.4 0.5-0.5 2.4-2.4 8.1-8.1 10.2-10.2 15.2-15.2 3.6-3.6
Gemini-1.5-Pro-002 $71.0$ $81.9$ 14.9-14.9 12.1-12.1 1.8-1.8 5.8-5.8 2.7-2.7 7.9-7.9 13.4-13.4 6.2-6.2
DeepSeek DeepSeek-V3 $69.1$ $82.9$ 11.5-11.5 6.6-6.6 +0.7+0.7 0.2-0.2 +1.2+1.2 +0.8+0.8 4.0-4.0 +2.6+2.6
LLaMA Series LLaMA-3.1-8B-Instruct $37.1$ $41.6$ 11.4-11.4 25.8-25.8 +1.3+1.3 +0.6+0.6 +1.1+1.1 1.5-1.5 2.7-2.7 10.4-10.4
LLaMA-3.1-70B-Instruct $62.1$ $66.4$ 17.6-17.6 11.8-11.8 +2.9+2.9 +3.0+3.0 +1.0+1.0 +1.2+1.2 10.8-10.8 6.6-6.6
LLaMA-3.1-405B-Instruct $66.8$ $75.0$ 10.2-10.2 6.8-6.8 +3.9+3.9 +2.8+2.8 +2.0+2.0 +1.9+1.9 5.3-5.3 4.2-4.2
LLaMA-3.3-70B-Instruct $63.3$ $76.5$ 18.5-18.5 11.9-11.9 +0.8+0.8 0.2-0.2 2.5-2.5 +0.5+0.5 13.8-13.8 8.5-8.5
Qwen Series Qwen2.5-7B-Instruct $38.8$ $51.4$ 17.2-17.2 25.9-25.9 +5.7+5.7 +1.7+1.7 +4.8+4.8 +3.9+3.9 3.7-3.7 7.4-7.4
Qwen2.5-14B-Instruct $50.4$ $60.8$ 19.3-19.3 9.1-9.1 +4.3+4.3 +6.6+6.6 +2.7+2.7 +3.2+3.2 4.3-4.3 0.8-0.8
Qwen2.5-32B-Instruct $63.5$ $74.1$ 16.5-16.5 15.1-15.1 +4.7+4.7 +0.4+0.4 +1.4+1.4 3.0-3.0 13.1-13.1 8.1-8.1
Qwen2.5-72B-Instruct $64.6$ $74.1$ 17.5-17.5 8.0-8.0 +1.7+1.7 +0.9+0.9 2.4-2.4 +1.5+1.5 16.5-16.5 5.4-5.4
Qwen2.5-Coder-32B-Instruct $74.2$ $78.4$ 21.2-21.2 16.1-16.1 0.8-0.8 0.1-0.1 1.8-1.8 +0.1+0.1 27.1-27.1 26.2-26.2
Average $\mathbf{-15.9}$ $\mathbf{-12.6}$ $\mathbf{+2.2}$ $\mathbf{+0.3}$ $\mathbf{+0.3}$ $\mathbf{-0.6}$ $\mathbf{-12.6}$ $\mathbf{-7.8}$
Appendix F Case Studies
F.1 Case Study 1: How is DeepSeek-V3 misled by PSC-ALL
@@
Code 19: PSC-ALL-perturbed example from Crux (sample 528) for output prediction.
@@ 18 Var_6 = Var_6 + Var_3
F.2 Case Study 2: How is GPT-4o misled by MCC
@@
Code 22: MCC-perturbed example from Crux (sample 1) for output prediction.
@@
F.3 Case Study 3: How is Qwen2.5-72B-Ins misled by MPS
This case (Code 24) highlights how MPS can subtly manipulate model reasoning. According to the CoT reasoning in Code 25, although the model initially performs all steps correctly and tracks how the occ is constructed, it misleads itself at the end by incorporating a print message into its logic. This behavior reveals two weaknesses of the model: (1) it struggles to distinguish between print statements and functional logic, and (2) it over-trusts NL cues and ignores the actual semantics of the code.
@@ 5. For each character ch in "URW rNB":
F.4 Case Study 4: How does MHC override Claude-3.5-Sonnet’s answer
@@ 5. Final value of ans is 4
F.5 Case Study 5: How does GPT-4o gain from reasoning shortcuts
@@
This function applies a sequence of operations to the input string text: (1) @ reverses the string, (2) ˜ converts it to uppercase, and (3) any other character removes itself from the end of the string if it matches the currently last character. LLMs are expected to provide input arguments that the function returns HI˜.
In the vanilla reasoning, as shown in Code 30, GPT-4o mentions, “Text is hi˜, ˜ makes it HI˜, @ reverses it to ˜IH ”, it wrongly assumes ˜IH is the final output because it applies the reversal after the uppercasing step, resulting in an incorrect answer.
In contrast, under the MCC perturbation as shown in Code 31, GPT-4o captures the embedded comments and states that “the comments indicate that most operations do not affect the final result.” Guided by these cues, the model seeks an input value for rules that avoids triggering any branches in the loop, effectively bypassing the actual code logic and getting the correct answer by leveraging reasoning shortcuts rather than truly understanding or executing the program. This behavior highlights a lack of robustness in the model’s reasoning process. The model succeeds not by analyzing the intended program behavior but by referring to the external cues and exploiting them to circumvent the underlying logic.
@@
Code 29: MCC-perturbed example from Crux (sample 623) for input prediction.
@@
F.6 Case Study 6: How does QwQ-32B collapse
In §3.4, we examine a failure case of QwQ-32B on LCB sample 259 (see Code 32), where the model collapses under MHC perturbation, generating over 32k tokens of reasoning. Figure 3 shows a quadratic growth of confusing tokens before the collapse. We manually segment the reasoning process into six stages (see Figure 3):
@@ 3 A = deque()
Appendix G Limitation
This study is subject to several limitations. Firstly, we mainly focus on algorithms and data structure questions in Python when constructing perturbations and evaluating model reliability using code reasoning. While our methods can be extended to realistic codebases and are directly transferable to other imperative languages, such as C/C++ and Java, they may not generalize to some languages like Prolog, ML, or Assembly because of the code semantics and structure. Secondly, although we evaluate 17 LLMs and 5 LRMs, many other latest closed-source models (e.g., GPT-4.5, o3, o4-Mini, Gemini-2.5, and Claude-3.7) and open-source models (e.g., Mixtral, Falcon, Vicuna) are not included due to budget constraints. Thirdly, we adopt Pass@1 as the primary evaluation metric. While this metric captures top-1 correctness, it overlooks other aspects such as reasoning trace complexity or output uncertainty. Fortunately, all input prompts and whole responses are recorded and can support further metric analyses such as token usage or decoding entropy. Lastly, our work does not offer definitive mitigation strategies. Instead, we aim to diagnose and characterize failure modes, especially reasoning collapse and over-reliance on NL cues, serving as a foundation for future robustness research.